### v1

@@ 3.4 Structural Consensus Constraint
In this work, we address these challenges by introducing a cyclic structural consensus constraint that bridges visual-language understanding in both forward and backward directions, as shown in Figure 5. Applied exclusively during the training stage, it serves as a powerful regularization mechanism without adding inference overhead. Specifically, we leverage our previously disentangled textual information to obtain representations for different semantic components through Equation. 2. This process yields three corresponding embeddings: static object embedding $\mathbf{e}_{so}\in\mathbb{R}^{1\times C}$, spatial motion embedding $\mathbf{e}_{sm}\in\mathbb{R}^{1\times C}$, and the global sentence representation $\mathbf{e}_{g}\in\mathbb{R}^{1\times C}$. Through cross-modal interaction, we derive the corresponding visual proxy features. The cross-attention mechanism facilitating this interaction is formulated as:
@@
where $\mathbf{e}$ represents the original text embedding and $\mathcal{D}_{t}(\cdot)$ denotes the transformer text reconstruction decoder.
@@
This reconstruction process enables the model to verify whether the tracked object sequence truly captures the semantic information specified in the original referring expression. To effectively validate the semantic alignment between tracked objects and their corresponding referring expressions, we need to quantify the consistency between original text embeddings and their reconstructed counterparts. However, establishing an appropriate consistency measure presents significant challenges. The point-wise consistency (Figure 5-B) constraints widely adopted in previous cross-modal tasks fall short in RMOT scenarios due to a fundamental limitation: they assume a bijective mapping between visual objects and textual descriptions. In real tracking scenarios, this assumption rarely holds as the same visual entity can be described through multiple valid linguistic expressions. Enforcing strict point-wise matching between original expression $\mathcal{E}$ and reconstructed embedding $\mathcal{E}^{\prime}$ would artificially constrain this natural language diversity, leading to suboptimal feature representations and degraded tracking performance.
@@
To address this limitation, we introduce a structural consensus constraint (Figure 5-C) that preserves semantic relationships while accommodating language variations. Instead of enforcing strict one-to-one correspondence between expressions, we maintain the relative structure of embedding spaces $\mathcal{E}$ and $\mathcal{E}^{\prime}$. The key insight is that while the same tracked object may have different valid descriptions, the semantic relationships between different objects should remain consistent. Our structural consensus constraint focuses on preserving two essential geometric properties across embedding spaces. Distance preservation ensures that relative distances between different referring expressions remain consistent in both original and reconstructed spaces, maintaining the integrity of semantic similarities. Simultaneously, angle preservation captures higher-order semantic structure by maintaining geometric relationships between expression triplets. When tracked objects correctly match their referring expressions, these structural properties remain well-preserved between $\mathcal{E}$ and $\mathcal{E}^{\prime}$, while mismatched trajectories manifest as structural inconsistencies. This approach allows our model to accommodate linguistic diversity while maintaining semantic consistency, significantly enhancing the robustness of multi-object tracking in language-guided scenarios. Formally, we define the structural consensus loss as:
@@ 4.1 Implementation Details
We adopt TransRMOT as our baseline architecture, which employs ResNet-50 [27] as the visual backbone and RoBERTa [41] as the text encoder. The text encoder’s parameters are kept frozen throughout the training process to maintain stable linguistic representations. The model is initialized with Deformable DETR [80] weights pre-trained on COCO dataset [39], except for the text encoder. We set the number of detect queries to $N=300$ to ensure sufficient coverage of potential objects. The architecture incorporates multi-scale feature learning with input resolutions ranging from 800 to 1536 pixels on the shortest side. The network is trained for 100 epochs on 2 NVIDIA RTX A5000 24GB GPUs with a batch size of 1 per GPU. We utilize AdamW [43] optimizer with a base learning rate of $1e^{-4}$, while the backbone’s learning rate is set to $1e^{-5}$. A learning rate decay of 0.1 is applied at the 50th epoch. For data augmentation, we employ random cropping and incorporate object erasing and inserting operations to enhance robustness to object entrances and exits, following [68]. The loss function is weighted with coefficients $\lambda_{cls}=2$, $\lambda_{L1}=5$, $\lambda_{giou}=2$, $\lambda_{ref}=2$, and $\lambda_{struct}=2$ to balance different training objectives.
During inference, CDRMT processes video sequences of arbitrary length in an online manner without requiring post-processing. At frame $t$, the model generates $N_{t}$ instance embeddings, each corresponding to either a true object or an empty prediction. We employ a two-stage filtering strategy: first, objects with confidence scores exceeding 0.7 are considered as valid detections; subsequently, the referring threshold $\beta_{ref}=0.5$ is applied to identify the final referred objects from these valid detections.
@@ 4.2.1 Datasets
4.2.2 Evaluation Metrics
Table 1: Comparison with state-of-the-art methods on Refer-KITTI-V1 dataset. $\uparrow$/ indicates higher scores are better. The best results are highlighted in bold. * means the result after frame correction strategy.
@@ 4.3.1 Evaluation on Refer-KITTI
In terms of overall tracking performance, CDRMT achieves the highest HOTA score of 49.35, surpassing our baseline TransRMOT* (46.56) by 6.0% and outperforming other competitive methods like iKUN* (48.84) and MLS-Track (49.05). This improvement demonstrates the effectiveness of our method in enhancing both detection and association capabilities. For detection-specific metrics, CDRMT demonstrates remarkable performance improvements over the TransRMOT* baseline. Specifically, our method achieves a DetA score of 40.34 (improved by 6.2% from 37.97) and a DetRe score of 54.54 (improved by 9.8% from 49.69). The DetPr score shows minimal difference (59.30 versus 60.10), indicating maintained precision while significantly improving recall capabilities. CDRMT shows competitive performance with an AssRe score of 64.70 (improved by 7.8% from TransRMOT*’s 60.02) and a slight AssPr score of 89.80 (improved by 0.1% from 89.67). These improvements validate our method’s enhanced capability in maintaining consistent object identities through language-guided association. Furthermore, CDRMT achieves state-of-the-art performance in spatial accuracy with a LocA score of 90.61, showing a slight but consistent improvement (0.3%) over TransRMOT*’s 90.33. This demonstrates our method’s ability to maintain precise object localization while enhancing overall tracking performance. When compared with other recent state-of-the-art methods, CDRMT shows comprehensive advantages. Notably, it outperforms the strong competitor iKUN* in multiple aspects, with improvements in HOTA (1.0%), DetA (12.9%), and association metrics. These consistent improvements across various metrics validate the effectiveness and robustness of our proposed approach in handling the challenging RMOT task.
@@ 4.3.4 Cross Comparison on Different Datasets
To comprehensively validate the effectiveness of our cognitive disentanglement method, we conducted extensive cross-method experiments by integrating CDRMT into two representative RMOT methods. As shown in Table 4, after incorporating our framework into TransRMOT [59], CDRMT demonstrates substantial improvements with relative gains of 6.0% HOTA, 6.2% DetA, and 5.6% AssA on Refer-KITTI, highlighting the benefits of disentangled semantic processing. More significantly, when enhancing TempRMOT [74] (denoted as CDRMT), our method exhibits consistent performance boosts with relative gains of 2.7% HOTA (from 50.68 to 52.04), 3.6% DetA (from 38.75 to 40.14), and 1.7% AssA (from 66.45 to 67.59). The improvements in both detection accuracy (DetA) and association quality (AssA) demonstrate that our cognitive-inspired design effectively enhances both spatial and temporal feature representations. Furthermore, these performance gains are consistently maintained on the more challenging Refer-KITTI-V2 dataset, where CDRMT achieves improvements of 1.8% HOTA, 2.8% DetA, and 0.8% AssA. These substantial and consistent improvements across different baseline methods and datasets validate that our cognitive-inspired disentangled processing of static and dynamic information can effectively enhance existing RMOT frameworks, leading to more robust and accurate tracking performance in complex scenarios.
@@ 4.4 Qualitative Results
As illustrated in the upper example of Figure 6, we observe distinct performance variations in response to the referring expression "women in the left". TransRMOT exhibits substantial false negatives (FN) across all frames. Specifically, in frame #$\#258$, it misses two female pedestrians in the object regions, and similar detection omissions persist in frame #$\#290$. While the method detects two objects in frame #$\#324$, it overlooks a crucial object near the left wall. Moreover, the erroneous attention to a male pedestrian on the right indicates insufficient or potentially misaligned integration of linguistic and visual features. Although both MOTRv2 and MGLT demonstrate comparable detection coverage, they still manifest detection insufficiencies across sequential frames. Notably, MOTRv2 exhibits substantially lower confidence scores compared to MGLT, which can be attributed to MGLT’s more sophisticated language-vision feature fusion mechanism. This performance gap underscores the importance of fine-grained multimodal feature integration in referring tracking tasks.
@@
Moreover, Figure 7 presents additional qualitative results comparing our CDRMT with state-of-the-art methods on the Refer-KITTI dataset, given the expression "black cars in the left". The visualization reveals distinct performance characteristics across different methods. TransRMOT exhibits significant detection instability, particularly manifesting false negatives in the initial frame and consistently low confidence scores throughout the sequence. While MGLT demonstrates improved object localization compared to TransRMOT, it suffers from imprecise bounding box regression in the first frame, where it incorrectly includes a white vehicle in the detection region. Furthermore, MGLT produces a false positive detection of a white vehicle in the second frame, indicating insufficient discrimination of color attributes specified in the referring expression. In contrast, our proposed CDRMT achieves superior performance in both object localization and attribute discrimination. The method demonstrates robust detection capabilities by accurately identifying black vehicles while maintaining precise bounding box boundaries across all frames. This enhanced performance can be attributed to our cognitive disentanglement framework, which effectively processes both static attributes ("black") and spatial information ("in the left") through separate pathways.
@@
Furthermore, we evaluated the performance of various methods in challenging scenes with the complex expression, as shown in Figure 8. The referring expression "left people who are walking" presents a significant challenge as it requires models to simultaneously understand object categories ("people"), spatial positioning ("left"), and motion states ("walking"). In frame #237, TransRMOT overlooks the rightmost pedestrian among the left group, failing to comprehend the motion component of the description. Meanwhile, MGLT incorrectly identifies stationary pedestrians on the left, demonstrating an inability to properly distinguish between static and dynamic objects. Similar issues persist in frame #268, where both TransRMOT and MGLT erroneously detect stationary pedestrians, indicating insufficient perception of motion information within the language description. In contrast, CDRMT accurately identifies only those pedestrians exhibiting walking behavior, effectively filtering out stationary individuals despite their spatial proximity. In frames #246 and #301, TransRMOT consistently misses walking pedestrians while incorrectly detecting pedestrians on the right side. While MGLT shows improved detection capabilities compared to TransRMOT, its performance still falls short of our CDRMT. These results validate the effectiveness of our cognitive disentanglement approach in handling complex language descriptions that combine both static object attributes and spatial motion information.

### v3

@@ 3.4 Structural Consistency Constraint
In this work, we address these challenges by introducing a structural consistency constraint that bridges visual-language understanding in both forward and backward directions, as shown in Figure 5. Applied exclusively during the training stage, it serves as a powerful regularization mechanism without adding inference overhead. Specifically, we leverage our previously disentangled textual information to obtain representations for different semantic components through Equation. 2. This process yields three corresponding embeddings: static object embedding $\mathbf{e}_{so}\in\mathbb{R}^{1\times C}$, spatial motion embedding $\mathbf{e}_{sm}\in\mathbb{R}^{1\times C}$, and the global sentence representation $\mathbf{e}_{g}\in\mathbb{R}^{1\times C}$. Through cross-modal interaction, we derive the corresponding visual proxy features. The cross-attention mechanism facilitating this interaction is formulated as:
@@
where $\mathcal{D}_{t}(\cdot)$ denotes the transformer text reconstruction decoder.
@@
This reconstruction process enables the model to verify whether the tracked object sequence truly captures the semantic information specified in the original referring expression. To effectively validate the semantic alignment between tracked objects and their corresponding referring expressions, we need to quantify the consistency between original text embeddings and their reconstructed counterparts. However, establishing an appropriate consistency measure presents significant challenges. The point-wise consistency (Figure 5-B) constraints widely adopted in previous cross-modal tasks fall short in RMOT scenarios due to a fundamental limitation: they assume a bijective mapping between visual objects and textual descriptions. From an information-theoretic standpoint, a point-wise matching constraint restricts the mutual information between visual and textual modalities. When considering the visual features $\mathbf{F}^{v}$ and text embeddings $\mathbf{e}$ as random variables, the point-wise matching effectively constrains their joint distribution $p(\mathbf{F}^{v},\mathbf{e})$ to be deterministic. However, in real-world RMOT scenarios, the optimal representation should maximize the mutual information $I(\mathbf{F}^{v},\mathbf{e})$ while accommodating the inherent many-to-many relationships between modalities:
$I(\mathbf{F}^{v},\mathbf{e})=\sum_{\mathbf{F}^{v},\mathbf{e}}p(\mathbf{F}^{v},\mathbf{e})\log\frac{p(\mathbf{F}^{v},\mathbf{e})}{p(\mathbf{F}^{v})p(\mathbf{e})},$ (16)
point-wise matching between the original expression $\mathcal{E}$ and the reconstructed embedding $\mathcal{E}^{\prime}$ would artificially constrain this information flow, leading to sub-optimal representations.
@@
To address this limitation, we introduce a structural consistency constraint (Figure 5-C) that preserves semantic relationships while accommodating language variations. Instead of enforcing strict one-to-one correspondence between expressions, we maintain the relative structure of embedding spaces $\mathcal{E}$ and $\mathcal{E}^{\prime}$. The key insight is that while the same tracked object may have different valid descriptions, the semantic relationships between different objects should remain consistent. Our structural consistency constraint focuses on preserving two essential geometric properties across embedding spaces. Distance preservation ensures that relative distances between different referring expressions remain consistent in both original and reconstructed spaces, maintaining the integrity of semantic similarities. Simultaneously, angle preservation captures higher-order semantic structure by maintaining geometric relationships between expression triplets. When tracked objects correctly match their referring expressions, these structural properties remain well-preserved between $\mathcal{E}$ and $\mathcal{E}^{\prime}$, while mismatched trajectories manifest as structural inconsistencies. This approach allows our model to accommodate linguistic diversity while maintaining semantic consistency, significantly enhancing the robustness of multi-object tracking in language-guided scenarios. Formally, we define the structural consistency loss as:
@@ 3.5 Loss Functions
Algorithm 1 Cognitive Disentanglement for Referring Multi-Object Tracking
0:  Video frames $\{\mathbf{F}_{t}\}_{t=1}^{T}$, language description $\mathcal{T}$
0:  Tracking results for referred objects
1:  $\mathbf{S}\leftarrow E_{l}(\mathcal{T})$, {Language encoding}
2:  $\{\mathcal{T}_{so},\mathcal{T}_{sm}\}\leftarrow\text{Dec}(\mathcal{T})$Semantic decoupling}
3:  for $t=1$ to $T$ do
4:     $\mathbf{V}_{t}\leftarrow E_{v}(\mathbf{F}_{t})$Visual encoding}
5:     // Bidirectional Interactive Fusion
6:     $\hat{\mathbf{V}}_{t},\hat{\mathbf{S}}\leftarrow\text{BIF}(\mathbf{V}_{t},\mathbf{S})$Cross-modal fusion}
7:     $\mathbf{F}^{v}_{t}\leftarrow\text{DeformEnc}(\hat{\mathbf{V}}_{t})\in\mathbb{R}^{HW\times d}$
8:     $\hat{\mathbf{f}}_{so},\hat{\mathbf{f}}_{sm}\leftarrow\text{ProcessTextFeatures}(\mathcal{T}_{so},\mathcal{T}_{sm},\hat{\mathbf{S}})$
9:     // Progressive Semantic-Decoupled Query Learning
10:     $\mathbf{Q}^{so}_{t}\leftarrow\text{QSI}(\mathbf{Q}_{t},\hat{\mathbf{f}}_{so})$What pathway}
11:     $\mathbf{Q}^{\prime}_{t}\leftarrow\text{DecLayers}_{0\rightarrow L/2}(\mathbf{Q}^{so}_{t},\mathbf{F}^{v}_{t})$
12:     $\hat{\mathbf{Q}}^{\prime}_{t}\leftarrow\text{DecLayers}_{L/2\rightarrow L}(\text{QSI}(\mathbf{Q}^{\prime}_{t},\hat{\mathbf{f}}_{sm}),\mathbf{F}^{v}_{t})$Where pathway}
13:     $\mathbf{B}_{t},\mathbf{C}_{t},\mathbf{R}_{t}\leftarrow\text{PredictionHeads}(\hat{\mathbf{Q}}^{\prime}_{t})$
14:     Association and tracking management if $t>1$
15:     Update queries for next frame
16:     if Training then
17:        // Structural Consensus Constraint
18:        $\mathbf{e}^{\prime}_{so},\mathbf{e}^{\prime}_{sm},\mathbf{e}^{\prime}_{g}\leftarrow\text{TextReconstruction}(\mathbf{F}^{v}_{t},\mathbf{F}^{so}_{t},\mathbf{F}^{sm}_{t})$
19:        $\mathcal{L}_{struct}\leftarrow\text{StructuralLoss}(\mathbf{e}_{so},\mathbf{e}_{sm},\mathbf{e}_{g},\mathbf{e}^{\prime}_{so},\mathbf{e}^{\prime}_{sm},\mathbf{e}^{\prime}_{g})$
20:        // RMOT losses
21:        $\mathcal{L}^{det}_{t}\leftarrow\sum^{N}_{i=1}[\lambda_{cls}\mathcal{L}_{cls}(\mathbf{c}^{det}_{t,i},\hat{\mathbf{c}}^{det}_{t,i})+\mathbb{1}_{\{\hat{\mathbf{c}}^{det}_{t,i}\neq\emptyset\}}(\mathcal{L}_{box}(\mathbf{b}^{det}_{t,i},\hat{\mathbf{b}}^{det}_{t,i})+\lambda_{ref}\mathcal{L}_{ref}(\mathbf{r}^{det}_{t,i},\hat{\mathbf{r}}^{det}_{t,i}))]$
22:        $\mathcal{L}^{tra}_{t}\leftarrow\sum^{N^{\prime}_{t-1}}_{i=1}[\lambda_{cls}\mathcal{L}_{cls}+\mathcal{L}_{box}+\lambda_{ref}\mathcal{L}_{ref}]$
23:        if $t=1$ then
24:           $\mathcal{L}_{t}\leftarrow\mathcal{L}^{det}_{t}+\lambda_{struct}\mathcal{L}_{struct}$
25:        else
26:           $\mathcal{L}_{t}\leftarrow\mathcal{L}^{tra}_{t}+\mathcal{L}^{det}_{t}+\lambda_{struct}\mathcal{L}_{struct}$
27:        end if
28:     end if
29:  end for
@@ 4.1 Implementation Details
We adopt TransRMOT as our baseline architecture, which employs ResNet-50 [25] as the visual backbone and RoBERTa [41] as the text encoder. The text encoder’s parameters are kept frozen throughout the training process to maintain stable linguistic representations. The model is initialized with Deformable DETR [83] weights pre-trained on the COCO dataset [39], except for the text encoder. We set the number of detection queries to $N=300$ to ensure sufficient coverage of potential objects. The network is trained for 100 epochs on 2 NVIDIA RTX A5000 24GB GPUs with a batch size of 1 per GPU. We utilize AdamW [43] optimizer with a base learning rate of $1e^{-4}$, while the backbone’s learning rate is set to $1e^{-5}$. A learning rate decay of 0.1 is applied at the 50th epoch. For data augmentation, we employ random cropping and incorporate object erasing and inserting operations to enhance robustness to object entrances and exits, following [69]. The loss function is weighted with coefficients $\lambda_{cls}=2$, $\lambda_{L1}=5$, $\lambda_{giou}=2$, $\lambda_{ref}=2$, and $\lambda_{struct}=2$ to balance different training objectives.
During inference, CDRMT processes video sequences of arbitrary length in an online manner without requiring post-processing. At frame $t$, the model generates $N_{t}$ instance embeddings, each corresponding to either a true object or an empty prediction. We employ a two-stage filtering strategy: first, objects with confidence scores exceeding 0.7 are considered as valid detections; subsequently, the referring threshold $\beta_{ref}=0.5$ is applied to identify the final referred objects from these valid detections. The entire pipeline of the method (CDRMT) is detailed in Algorithm 1.
@@ 4.2.1 Datasets
Table 1: Comparison with state-of-the-art methods on Refer-KITTI dataset. $\uparrow$/ indicates higher scores are better. The best results are highlighted in bold. * means the result after frame correction strategy.
@@ 4.3.1 Evaluation on Refer-KITTI
In terms of overall tracking performance, CDRMT achieves the highest HOTA score of 49.35, surpassing our baseline TransRMOT* (46.56) by 6.0% and outperforming other competitive methods like iKUN* (48.84) and MLS-Track (49.05). This improvement demonstrates the effectiveness of our method in enhancing both detection and association capabilities. For detection-specific metrics, CDRMT demonstrates remarkable performance improvements over the TransRMOT* baseline. Specifically, our method achieves a DetA score of 40.34 (improved by 6.2% from 37.97) and a DetRe score of 54.54 (improved by 9.8% from 49.69). The DetPr score shows minimal difference (59.30 versus 60.10). CDRMT shows competitive performance with an AssRe score of 64.70 (improved by 7.8% from TransRMOT*’s 60.02) and a slight AssPr score of 89.80 (improved by 0.1% from 89.67). These improvements validate our method’s enhanced capability in maintaining consistent object identities through language-guided association. Furthermore, CDRMT achieves state-of-the-art performance in spatial accuracy with a LocA score of 90.61, showing a slight but consistent improvement (0.3%) over TransRMOT*’s 90.33. This demonstrates our method’s ability to maintain precise object localization while enhancing overall tracking performance. When compared with other recent state-of-the-art methods, CDRMT shows comprehensive advantages. Notably, it outperforms the strong competitor iKUN* in multiple aspects, with improvements in HOTA (1.0%), DetA (12.9%), and association metrics. These consistent improvements across various metrics validate the effectiveness and robustness of our proposed approach in handling the challenging RMOT task.
@@ 4.3.4 Cross Comparison on Different Datasets
To comprehensively validate the effectiveness of our cognitive disentanglement method, we conducted extensive cross-method experiments by integrating CDRMT into two representative RMOT methods. As shown in Table 4, after incorporating our framework into TransRMOT [60], CDRMT demonstrates substantial improvements with relative gains of 6.0% HOTA, 6.2% DetA, and 5.6% AssA on Refer-KITTI, highlighting the benefits of disentangled semantic processing. More significantly, when enhancing TempRMOT [77] (denoted as CDRMT), our method still yields consistent performance boosts, albeit with comparatively smaller relative gains of 2.7% HOTA (from 50.68 to 52.04), 3.6% DetA (from 38.75 to 40.14), and 1.7% AssA (from 66.45 to 67.59). This performance differential between TransRMOT and TempRMOT integration stems from several factors. TempRMOT begins from a substantially higher baseline performance, typically leading to diminishing returns as models approach higher performance levels. Additionally, TempRMOT’s sophisticated temporal query memory mechanism likely captures some motion information targeted by our spatial-motion pathway, creating partial representational overlap. Our primary goal in the transfer study was to demonstrate the general applicability of our approach rather than to maximize performance gains through extensive architectural tuning. Nevertheless, the improvements in both detection accuracy (DetA) and association quality (AssA) demonstrate that our cognitive-inspired design effectively enhances both spatial and temporal feature representations. Furthermore, these performance gains are consistently maintained on the more challenging Refer-KITTI-V2 dataset, where CDRMT achieves improvements of 1.8% HOTA, 2.8% DetA, and 0.8% AssA. These substantial and consistent improvements across different baseline methods and datasets validate that our cognitive-inspired disentangled processing of static and dynamic information can effectively enhance existing RMOT frameworks, leading to more robust and accurate tracking performance in complex scenarios.
@@ 4.4 Qualitative Results
Figure 9: Qualitative comparison of referring tracking results with baseline methods on expressions without explicit spatial information.
As illustrated in the upper example of Figure 6, we observe distinct performance variations in response to the referring expression "women in the left". TransRMOT exhibits substantial false negatives (FN) across all frames. Specifically, in frame #$\#258$, it misses two female pedestrians in the object regions, and similar detection omissions persist in frame #$\#290$. While the method detects two objects in frame #$\#324$, it overlooks a crucial object near the left wall. Moreover, the erroneous attention of MOTRv2 to a male pedestrian on the right indicates insufficient or potentially misaligned integration of linguistic and visual features. Although both MOTRv2 and MGLT demonstrate comparable detection coverage, they still manifest detection insufficiencies across sequential frames. Notably, MOTRv2 exhibits substantially lower confidence scores compared to MGLT, which can be attributed to MGLT’s more sophisticated language-vision feature fusion mechanism. This performance gap underscores the importance of fine-grained multimodal feature integration in referring tracking tasks.
@@
Moreover, Figure 7 presents additional qualitative results comparing our CDRMT with state-of-the-art methods on the Refer-KITTI dataset, given the expression "black cars in the left". The visualization reveals distinct performance characteristics across different methods. TransRMOT exhibits significant detection instability, particularly manifesting false negatives in the initial frame and consistently low confidence scores throughout the sequence. While MGLT demonstrates improved object localization compared to TransRMOT, it suffers from imprecise bounding box regression in the first frame, where it incorrectly includes a white vehicle in the detection region. Furthermore, MGLT produces a false positive detection of a white vehicle in the second frame, indicating insufficient discrimination of color attributes specified in the referring expression. In contrast, our proposed CDRMT achieves superior performance in both object localization and attribute discrimination. The method demonstrates robust detection capabilities by accurately identifying black vehicles while maintaining precise bounding box boundaries. This enhanced performance can be attributed to our cognitive disentanglement framework, which effectively processes both static attributes ("black") and spatial information ("in the left") through separate pathways.
@@
Furthermore, we evaluated the performance of various methods in challenging scenes with complex expressions, as shown in Figure 8. The referring expression "left people who are walking" presents a significant challenge as it requires models to simultaneously understand object categories ("people"), spatial positioning ("left"), and motion states ("walking"). In frame #237, TransRMOT overlooks the rightmost pedestrian among the left group, failing to comprehend the motion component of the description. Meanwhile, MGLT incorrectly identifies stationary pedestrians on the left, demonstrating an inability to properly distinguish between static and dynamic objects. Similar issues persist in frame #268, where both TransRMOT and MGLT erroneously detect stationary pedestrians, indicating insufficient perception of motion information within the language description. In contrast, CDRMT accurately identifies only those pedestrians exhibiting walking behavior, effectively filtering out stationary individuals despite their spatial proximity. In frames #246 and #301, TransRMOT consistently misses walking pedestrians while incorrectly detecting pedestrians on the right side. While MGLT shows improved detection capabilities compared to TransRMOT, its performance still falls short of our CDRMT. These results validate the effectiveness of our cognitive disentanglement approach in handling complex language descriptions that combine both static object attributes and spatial motion information.
As illustrated in Figure 9, we evaluate our model against baselines on expressions lacking explicit spatial localization terms. For the expression "vehicles in light color," which primarily tests static attribute recognition capability, TransRMOT exhibits significant detection instability. It fails to identify the target objects in frames #84 and #193, and even when detection occurs in frame #231, it shows substantially lower confidence (0.89) compared to our method (0.98). In contrast, CDRMT consistently identifies all light-colored vehicles across all frames with high confidence scores. It demonstrates the effectiveness of our enhanced static attribute processing through the "what" pathway, even in the absence of spatial indicators. The expression "moving vehicles," which requires understanding of motion dynamics without explicit positional cues, the limitations of baseline methods become more pronounced. TransRMOT incorrectly identifies stationary vehicles as targets across frames #186, #190, and #275, failing to distinguish motion attributes. Our CDRMT framework, however, correctly recognizes that motion terms like "moving" should be processed through the "where" pathway despite lacking traditional spatial prepositions. By effectively decoupling this motion information, CDRMT successfully identifies only the vehicles in motion while filtering out stationary vehicles that don’t match the specified criteria. This demonstrates our model’s superior capability in handling complex semantic relationships that combine both static object recognition and dynamic state understanding.
Table 5: Ablation studies of different components in CDRMT. In the table, ‘w.’ indicates ‘with’ (component used) and ‘w/o.’ indicates ‘without’ (component not used). The best results are in bold.
Components Metrics
[…]
46.56 37.97 57.33 49.69 60.10 60.02 89.67 90.33
46.96 38.65 57.11 55.75 54.64 61.35 88.73 90.43
48.36 39.97 58.69 54.86 58.07 62.79 89.68 90.32
w/o. ${Q}^{\prime}_{t}$ 48.02 39.45 58.62 55.46 56.41 63.09 88.77 90.52
w/o. $\hat{Q}^{\prime}_{t}$ 47.89 39.63 58.04 54.66 57.64 61.94 89.77 90.60
48.67 40.05 59.31 55.87 57.89 63.13 88.57 90.51
47.71 39.13 58.34 53.97 55.43 62.61 88.75 90.23
49.06 40.11 60.13 54.97 58.72 62.95 89.24 90.57
49.35 40.34 60.56 54.54 59.30 64.70 89.80 90.61